# Randomized HyperSteiner: A Stochastic Delaunay Triangulation Heuristic for the Hyperbolic Steiner Minimal Tree

### v1

@@ 1 Introduction
The principle of maximum parsimony, in line with Occam’s razor, states that among competing explanations consistent with the evidence, the simplest should be preferred. In phylogenetics, for example, this principle manifests through tree-based models, where minimality of length provides a natural notion of parsimony (Fitch,, 1971). Steiner Minimal Trees (SMTs) formalize this objective by seeking the shortest tree connecting a set of points, potentially augmented with additional Steiner points. In contrast to the Minimum Spanning Tree (MST), which only uses the input terminals, SMTs may reduce total length substantially. This makes SMTs a particularly compelling model when inferring hierarchical structures from data, where the ground truth tree is typically unknown, while tree length is an interpretable and measurable quantity that allows systematic comparison across methods.
@@ 2 Related work
Euclidean Delaunay-Based Heuristic SMTs. Heuristic methods for Euclidean Steiner Minimal Trees (SMTs) commonly exploit geometric structures, particularly Delaunay triangulations (DT), to identify promising local configurations for optimization (Zachariasen and Winter,, 1999). These algorithms typically follow a three-phase structure: Expansion selects triangles from the DT as candidates for local 3-point Steiner problems (i.e., Fermat points); Heuristic Construction assembles local solutions into a global tree; and Exploration Policy determines refinement strategies. The seminal Smith-Lee-Liebman (SLL) algorithm (Smith et al.,, 1981) exemplifies this framework through deterministic MST-based expansion. SLL selects Delaunay triangles containing two MST edges as candidates for Steiner point insertion. When all internal angles are below 120°, the triangle is replaced by a its Fermat point, minimizing total edge length while ensuring 120° edge angles (Gilbert and Pollak,, 1968). This yields locally-optimal full Steiner topologies with high computational efficiency. However, deterministic expansion strategies suffer from limited topological exploration, often becoming trapped in local minima for clustered or symmetric point configurations (Laarhoven and Ohlmann,, 2011). To address this limitation, researchers have introduced stochasticity at multiple algorithmic levels. Beasley and Goffinet, (1994) pioneered this direction with two key innovations: iterative expansion that recomputes the DT after each Steiner point addition, enabling dynamic geometric adaptation; and simulated annealing as stochastic exploration policy. Beasley’s method examines all DT triangles at each expansion step, ensuring comprehensive local configuration analysis. Laarhoven and Ohlmann, (2011) further advanced this approach by demonstrating that stochastic triangle sampling achieves comparable results with significantly reduced computational cost. Instead of exhaustively processing all triangles, their method randomly samples Delaunay simplices during expansion, improving solution quality through broader topological exploration while maintaining efficiency. Similarly, our proposed approach employs stochastic exploration to overcome the limitations of deterministic hyperbolic methods such as García-Castellanos et al., (2025)
@@ 3.1 The Klein-Beltrami Model
where $\langle\cdot,\cdot\rangle$ denotes the Lorentzian Inner Product in homogeneous coordinates, $I_{n}$ the identity matrix, and $\otimes$ the tensor product. The corresponding geodesic distance between two points $x,y\in\mathbb{K}^{n}$ is given by:
@@ 4 Method
 Initialize $n\leftarrow 1$, $S\leftarrow\emptyset$, $T^{*}\leftarrow\text{MST}(P)$.
@@ 4.1 Hyperbolic Delaunay Triangulation
Given $P\subseteq\mathbb{K}^{n}$, there exists an explicit set $S\subseteq\mathbb{R}^{n}$ and weights $\{r_{s}\}_{s\in S}$ such that the hyperbolic Voronoi cells of $P$ correspond to restrictions to $\mathbb{K}^{n}$ of power cells of $S$.
@@ 4.2 Local FST Solutions via Fermat Points
The resulting nonlinear system can be solved via numerical methods, such as the hybrid algorithm proposed by Powell, (1970), which extends Newton’s method to higher dimensions for root finding.
@@ 4.3 Global Refinement via Riemannian GD
where $\exp$ denotes the Riemannian exponential map and $\nabla L$ the Riemannian gradient of total tree length. This ensures descent along geodesics, extending Euclidean GD to hyperbolic space. Convergence properties mirror the Euclidean case presented in Gilbert and Pollak, (1968). Indeed, we show the following uniqueness result:
@@ 5 Experiments
We empirically evaluate Randomized HyperSteiner (RHS) for constructing hyperbolic Steiner minimal trees with two objectives: (i) quantify tree quality via reduction over the Minimum Spanning Tree (RED, %), and (ii) benchmark against standard methods—Minimum Spanning Tree (MST), Neighbour Joining (NJ) (Saitou and Nei,, 1987), and the deterministic HyperSteiner (HS) heuristic (García-Castellanos et al.,, 2025). Experiments were conducted on Google Cloud (n2d CPUs, 1 core, 128GB RAM). Additional details appear in Appendix D.
@@ 5.1 Performance Across Data Distributions
Results. Figure 3 demonstrates that RHS consistently outperforms all baselines. In the centered Gaussian regime (Figure 3(a)), the dense central distribution yields Euclidean behavior as $|P|$ increases due to hyperbolic space’s local flatness, producing modest RHS improvements over HS ($\sim$1%) comparable to 2D Euclidean stochastic DT-based heuristics (Laarhoven and Ohlmann,, 2011). Randomization becomes critical in the boundary mixture (Figure 3(b)): RHS maintains substantial reductions as sample size grows, while HS provides only moderate gains and NJ deteriorates to negative performance. On real biological data (Figure 3(c)), RHS sustains stable improvements over HS, whereas NJ exhibits inconsistent behavior. Notably, despite being a common computational biology baseline, NJ fails to respect the principle of maximum parsimony. The marked performance advantage for boundary-concentrated clusters (Figure 3(b)) motivates a deeper investigation of the relationship between boundary proximity and tree-length reduction.
@@ 5.2 Near Boundary Performance Analysis
The substantial improvements for boundary-concentrated clusters raise two fundamental questions: (i) how closely can heuristic methods approach the theoretical upper bound on reduction, and (ii) at what radial distance do these gains materialize? We address both through controlled experiments that systematically vary boundary proximity. Additional analysis appears in Appendix E.
@@ 6 Discussion and Future Work
We introduced Randomized HyperSteiner (RHS), a stochastic Delaunay triangulation heuristic that overcomes the fundamental limitation of deterministic approaches to hyperbolic Steiner Minimal Tree construction. By incorporating randomness into the expansion process and combining it with Riemannian gradient descent optimization, RHS avoids the local optima that trap existing methods.
@@
Despite these advances, our results reveal that we have not yet achieved the full potential of hyperbolic geometry for tree minimization. A gap of approximately 7% remains between our best empirical results and the theoretical upper bound. A promising direction for future work involves developing Polynomial-Time Approximation Schemes (PTAS) using quadtrees instead of Delaunay triangulations (Arora et al.,, 1998; Arora,, 1998; Mitchell,, 1999). PTAS approaches would allow users to explicitly control the performance-time trade-off by adjusting the approximation parameter. While the construction of quadtrees in hyperbolic space has been shown to be theoretically feasible (Kisfaludi-Bak and van Wordragen,, 2025), the practical implementation of these polynomial-time approximation algorithms in hyperbolic geometry remains an open challenge that could finally bridge the gap to optimal performance.
@@ E.1 Performance Across Data Distributions
NJ HS RHS (ours)
$|P|$ RED $(\uparrow)$ CPU RED $(\uparrow)$ CPU RED $(\uparrow)$ CPU
10 $1.46_{\pm 1.12}$ $5.50_{\pm 0.03}$ $2.72_{\pm 0.97}$ $0.03_{\pm 0.01}$ $3.17_{\pm 1.59}$ $4.60_{\pm 3.40}$
20 $2.60_{\pm 3.00}$ $5.68_{\pm 0.04}$ $1.83_{\pm 0.70}$ $0.06_{\pm 0.01}$ $3.15_{\pm 0.56}$ $9.87_{\pm 7.00}$
30 $1.70_{\pm 2.31}$ $5.86_{\pm 0.03}$ $2.12_{\pm 0.60}$ $0.10_{\pm 0.02}$ $3.03_{\pm 1.09}$ $12.65_{\pm 6.94}$
40 $2.62_{\pm 1.35}$ $5.95_{\pm 0.04}$ $2.17_{\pm 0.71}$ $0.13_{\pm 0.02}$ $3.30_{\pm 0.75}$ $17.59_{\pm 6.97}$
50 $4.37_{\pm 1.87}$ $6.14_{\pm 0.04}$ $2.26_{\pm 0.43}$ $0.17_{\pm 0.02}$ $3.08_{\pm 0.34}$ $26.21_{\pm 8.67}$
60 $3.38_{\pm 1.09}$ $6.29_{\pm 0.03}$ $2.56_{\pm 0.57}$ $0.20_{\pm 0.02}$ $2.72_{\pm 0.48}$ $26.95_{\pm 8.40}$
70 $3.28_{\pm 1.48}$ $6.51_{\pm 0.04}$ $2.45_{\pm 0.54}$ $0.22_{\pm 0.01}$ $2.68_{\pm 0.66}$ $31.09_{\pm 11.84}$
80 $3.89_{\pm 1.79}$ $6.77_{\pm 0.03}$ $2.43_{\pm 0.53}$ $0.26_{\pm 0.02}$ $3.07_{\pm 0.39}$ $38.55_{\pm 8.45}$
90 $3.46_{\pm 1.65}$ $7.01_{\pm 0.04}$ $2.72_{\pm 0.50}$ $0.30_{\pm 0.02}$ $2.87_{\pm 0.57}$ $66.07_{\pm 30.15}$
100 $3.77_{\pm 1.75}$ $7.24_{\pm 0.03}$ $2.45_{\pm 0.35}$ $0.33_{\pm 0.02}$ $2.84_{\pm 0.47}$ $70.31_{\pm 14.92}$
150 $4.17_{\pm 1.10}$ $8.46_{\pm 0.03}$ $2.27_{\pm 0.20}$ $0.48_{\pm 0.02}$ $2.48_{\pm 0.22}$ $139.87_{\pm 49.03}$
200 $5.01_{\pm 1.40}$ $11.09_{\pm 0.06}$ $2.43_{\pm 0.26}$ $0.66_{\pm 0.04}$ $2.64_{\pm 0.23}$ $341.42_{\pm 96.02}$
@@
Table 1: Sampling $|P|$ points from a centered hyperbolic Gaussian $\mathcal{G}(0,0.5)$. RED: reduction over MST (%). CPU: total CPU time (sec.).
NJ HS RHS (ours)
$|P|$ RED $(\uparrow)$ CPU RED $(\uparrow)$ CPU RED $(\uparrow)$ CPU
10 $40.49_{\pm 0.03}$ $5.52_{\pm 0.04}$ $17.75_{\pm 4.57}$ $0.02_{\pm 0.01}$ $28.33_{\pm 18.54}$ $25.70_{\pm 13.30}$
20 $40.10_{\pm 0.06}$ $5.73_{\pm 0.05}$ $14.35_{\pm 4.61}$ $0.05_{\pm 0.02}$ $40.01_{\pm 0.09}$ $61.94_{\pm 16.77}$
30 $39.93_{\pm 0.08}$ $5.91_{\pm 0.04}$ $14.74_{\pm 5.75}$ $0.08_{\pm 0.06}$ $39.84_{\pm 0.13}$ $95.05_{\pm 27.50}$
40 $39.70_{\pm 0.09}$ $6.06_{\pm 0.03}$ $16.63_{\pm 4.46}$ $0.09_{\pm 0.03}$ $39.71_{\pm 0.12}$ $119.86_{\pm 53.15}$
50 $39.53_{\pm 0.06}$ $6.00_{\pm 0.02}$ $17.88_{\pm 5.29}$ $0.10_{\pm 0.04}$ $39.52_{\pm 0.05}$ $116.38_{\pm 45.92}$
60 $39.22_{\pm 0.10}$ $6.37_{\pm 0.04}$ $14.95_{\pm 4.60}$ $0.11_{\pm 0.02}$ $39.36_{\pm 0.11}$ $87.99_{\pm 9.33}$
70 $39.01_{\pm 0.16}$ $6.55_{\pm 0.05}$ $14.42_{\pm 4.27}$ $0.11_{\pm 0.03}$ $39.26_{\pm 0.05}$ $118.26_{\pm 24.23}$
80 $38.55_{\pm 0.11}$ $6.77_{\pm 0.04}$ $14.47_{\pm 4.29}$ $0.12_{\pm 0.03}$ $39.11_{\pm 0.14}$ $111.56_{\pm 12.72}$
90 $38.18_{\pm 0.18}$ $6.96_{\pm 0.06}$ $15.85_{\pm 3.70}$ $0.14_{\pm 0.03}$ $39.03_{\pm 0.08}$ $171.84_{\pm 59.68}$
100 $37.50_{\pm 0.28}$ $7.26_{\pm 0.07}$ $13.91_{\pm 4.99}$ $0.13_{\pm 0.04}$ $38.94_{\pm 0.12}$ $183.57_{\pm 60.30}$
150 $24.02_{\pm 1.24}$ $8.48_{\pm 0.05}$ $13.22_{\pm 6.21}$ $0.16_{\pm 0.02}$ $38.47_{\pm 0.10}$ $205.34_{\pm 52.63}$
200 $8.08_{\pm 1.09}$ $10.51_{\pm 0.06}$ $13.62_{\pm 4.43}$ $0.18_{\pm 0.03}$ $38.17_{\pm 0.17}$ $265.46_{\pm 80.59}$
Table 2: Sampling $|P|$ points from a mixture of $10$ hyperbolic Gaussians $\mathcal{G}(\mu_{10,k}(1-10^{-10}),0.1)$, $k\in\{1,\cdots,10\}$. RED: reduction over MST (%). CPU: total CPU time (sec.).

### v2

@@ 1 Introduction
The principle of maximum parsimony, in line with Occam’s razor, states that among competing explanations consistent with the evidence, the simplest should be preferred. In phylogenetics, for example, this principle manifests through tree-based models, where minimality of length provides a natural notion of parsimony (Fitch,, 1971). Steiner Minimal Trees (SMTs) formalize this objective by seeking the shortest tree connecting a set of points, potentially augmented with additional Steiner points. In contrast to the Minimum Spanning Tree (MST), which only uses the input terminals, SMTs may reduce total length substantially. This makes SMTs a particularly compelling model when inferring hierarchical structures from data, where the ground truth tree is typically unknown, while tree length is an interpretable and measurable quantity that allows systematic comparison across methods. From a broader perspective, our main contribution lies in non-Euclidean computational geometry, with biological hierarchical inference serving only as one potential application of the proposed framework.
11footnotetext: Equal contribution. E-mails: <<medbouhi@kth.se>>, <<a.garciacastellanos@uva.nl>>.
@@ 2 Related work
Euclidean Delaunay-Based Heuristic SMTs. Heuristic methods for Euclidean Steiner Minimal Trees (SMTs) commonly exploit geometric structures, particularly Delaunay triangulations (DT), to identify promising local configurations for optimization (Zachariasen and Winter,, 1999). These algorithms typically follow a three-phase structure: Expansion selects triangles from the DT as candidates for local 3-point Steiner problems (i.e., Fermat points); Heuristic Construction assembles local solutions into a global tree; and Exploration Policy determines refinement strategies. The seminal Smith-Lee-Liebman (SLL) algorithm (Smith et al.,, 1981) exemplifies this framework through deterministic MST-based expansion. SLL selects Delaunay triangles containing two MST edges as candidates for Steiner point insertion. When all internal angles are below 120°, the triangle is replaced by its Fermat point, minimizing total edge length while ensuring 120° edge angles (Gilbert and Pollak,, 1968). This yields locally-optimal full Steiner topologies with high computational efficiency. However, deterministic expansion strategies suffer from limited topological exploration, often becoming trapped in local minima for clustered or symmetric point configurations (Laarhoven and Ohlmann,, 2011). To address this limitation, researchers have introduced stochasticity at multiple algorithmic levels. Beasley and Goffinet, (1994) pioneered this direction with two key innovations: iterative expansion that recomputes the DT after each Steiner point addition, enabling dynamic geometric adaptation; and simulated annealing as stochastic exploration policy. Beasley’s method examines all DT triangles at each expansion step, ensuring comprehensive local configuration analysis. Laarhoven and Ohlmann, (2011) further advanced this approach by demonstrating that stochastic triangle sampling achieves comparable results with significantly reduced computational cost. Instead of exhaustively processing all triangles, their method randomly samples Delaunay simplices during expansion, improving solution quality through broader topological exploration while maintaining efficiency. Similarly, our proposed approach employs stochastic exploration to overcome the limitations of deterministic hyperbolic methods such as García-Castellanos et al., (2025)
@@ 3.1 The Klein-Beltrami Model
where $\langle\cdot,\cdot\rangle:=-1+\langle\cdot,\cdot\rangle_{E}$ denotes the Lorentzian Inner Product in homogeneous coordinates (with $\langle\cdot,\cdot\rangle_{E}$ being the standard Euclidean inner product), $I_{n}$ the identity matrix, and $\otimes$ the tensor product. The corresponding geodesic distance between two points $x,y\in\mathbb{K}^{n}$ is given by:
@@ 4 Method
 Initialize $n\leftarrow 1$, $\leftarrow\emptyset, SS^{*}\leftarrow\emptyset$, $T^{*}\leftarrow\text{MST}(P)$.
@@ 4.1 Hyperbolic Delaunay Triangulation
Given $P\subseteq\mathbb{K}^{n}$, there exists an explicit set $\widetilde{P}\subseteq\mathbb{R}^{n}$ and weights $\{r_{\widetilde{p}}\}_{\widetilde{p}\in\widetilde{P}}$ such that the hyperbolic Voronoi cells of $P$ correspond to restrictions to $\mathbb{K}^{n}$ of power cells of $\widetilde{P}$.
Specifically, the set $\widetilde{P}$ (for computing power cells) and weights $\{r_{\widetilde{p}}\}_{\widetilde{p}\in\widetilde{P}}$ are derived from points $p\in P$ as follows:
$\widetilde{p}=\frac{p}{2\sqrt{1-\|p\|^{2}}},\hskip 30.00005ptr_{\widetilde{p}}^{2}=\frac{\|p\|^{2}}{4(1-\|p\|^{2})}-\frac{1}{\sqrt{1-\|p\|^{2}}}.$ (5)
@@ 4.2 Local FST Solutions via Fermat Points
The resulting nonlinear system can be solved numerically in $\mathcal{O}(1)$ time (see Appendix E.4 for details), for example via the hybrid algorithm proposed by Powell, (1970), which extends Newton’s method to higher dimensions for root finding.
@@ 4.3 Global Refinement via Riemannian GD
where $\exp$ denotes the Riemannian exponential map (applied to each Steiner point) and $\nabla L$ the Riemannian gradient of total tree length. This ensures descent along geodesics, extending Euclidean GD to hyperbolic space. Convergence properties mirror the Euclidean case presented in Gilbert and Pollak, (1968). Indeed, we show the following uniqueness result:
@@ 5 Experiments
We empirically evaluate Randomized HyperSteiner (RHS) for constructing hyperbolic Steiner minimal trees with two objectives: (i) quantify tree quality via reduction over the Minimum Spanning Tree (RED, %), and (ii) benchmark against standard methods—Minimum Spanning Tree (MST), Neighbour Joining (NJ) (Saitou and Nei,, 1987), and the deterministic HyperSteiner (HS) heuristic (García-Castellanos et al.,, 2025). Experiments were conducted on Google Cloud (n2d CPUs, 1 core, 128GB RAM). Source code is available in https://github.com/AGarciaCast/RandomizedHyperSteiner; additional details appear in Appendix D.
@@ 5.1 Performance Across Data Distributions
Results.
Figure 3 demonstrates that RHS consistently outperforms all baselines. In the centered Gaussian regime (Figure 3(a)), the dense central distribution yields Euclidean behavior as $|P|$ increases due to hyperbolic space’s local flatness, producing modest RHS improvements over HS ($\sim$1%) comparable to 2D Euclidean stochastic DT-based heuristics (Laarhoven and Ohlmann,, 2011). Randomization becomes critical in the boundary mixture (Figure 3(b)): RHS maintains substantial reductions as sample size grows, while HS provides only moderate gains and NJ deteriorates to negative performance. On real biological data (Figure 3(c)), RHS sustains stable improvements over HS, whereas NJ exhibits inconsistent behavior. Notably, despite being a common computational biology baseline, NJ fails to respect the principle of maximum parsimony. The marked performance advantage for boundary-concentrated clusters (Figure 3(b)) motivates a deeper investigation of the relationship between boundary proximity and tree-length reduction.
@@ 5.2 Near Boundary Performance Analysis
The substantial improvements for boundary-concentrated clusters raise two fundamental questions: (i) how closely can heuristic methods approach the theoretical upper bound on reduction, and (ii) at what radial distance do these gains materialize? We address both through controlled experiments that systematically vary boundary proximity.
@@
6 Theoretical Limitations of Stochastic Topology Search
Our method shares a key limitation with prior Euclidean stochastic Steiner tree heuristics (Beasley and Goffinet,, 1994; Laarhoven and Ohlmann,, 2011; Yang,, 2006): for general point clouds, it does not provide a theoretical guarantee of finding the globally optimal topology. While Theorem 4.3 shows that, for a fixed topology, the Steiner point configuration is uniquely determined, and thus the global refinement step showcased in Section 4.3 is well posed, the topology itself remains unknown and must be explored heuristically.
HS: $|P|=5$
RHS: $|P|=5$
HS: $|P|=6$
RHS: $|P|=6$
HS: $|P|=7$
RHS: $|P|=7$
[…]
Accordingly, RHS relies on stochastic exploration over candidate topologies while retaining the best solution found so far. This makes standard convergence analysis infeasible. In particular, changes in topology can produce discontinuous changes in tree length, which makes it difficult to obtain a straightforward characterization of convergence or theoretical computational complexity, as in other non-deterministic Steiner tree methods (Beasley and Goffinet,, 1994; Laarhoven and Ohlmann,, 2011).
Despite these theoretical limitations, the results in Section 5 and Appendix E show that the proposed pipeline consistently improves over existing baselines across a range of synthetic and real hyperbolic datasets.
@@ 7 Discussion and Future Work
We introduced Randomized HyperSteiner (RHS), a stochastic Delaunay triangulation heuristic that overcomes the fundamental limitation of deterministic approaches to constructing hyperbolic Steiner Minimal Trees. By incorporating randomness into the expansion process and combining it with Riemannian gradient descent optimization, RHS avoids the local optima that trap existing methods. Moreover, as shown in Figure 6, the myopic behavior of the original HyperSteiner prevents it from generating complex hierarchical trees, whereas our proposed RHS, thanks to our iterative exploration procedure, is able to construct richer hierarchical structures with shorter total tree length.
@@
Despite these advances, our results reveal that we have not yet achieved the full potential of hyperbolic geometry for tree length minimization. A gap of approximately 7% remains between our best empirical results and the theoretical upper bound. A promising direction for future work involves developing Polynomial-Time Approximation Schemes (PTAS) using quadtrees instead of Delaunay triangulations (Arora et al.,, 1998; Arora,, 1998; Mitchell,, 1999). PTAS approaches would allow users to explicitly control the performance-time trade-off by adjusting the approximation parameter. While the construction of quadtrees in hyperbolic space has been shown to be theoretically feasible (Kisfaludi-Bak and van Wordragen,, 2025), the practical implementation of these polynomial-time approximation algorithms in hyperbolic geometry remains an open challenge that could finally bridge the gap to optimal performance.
@@ A.1 Expansion Phase
Detailed ablations regarding the insertion probabilities and the expansion scheduler can be found in Section E.5 and Section E.6.
@@ D.1 Hyperbolic SMT Heuristic Methods
Although optimal learning rates are dataset-dependent, extensive hyperparameter tuning shows that $\eta=1\times 10^{-2}$ performs consistently well across all datasets. As noted in Appendix A.3, more advanced optimization schemes (e.g., Armijo backtracking (Absil et al.,, 2008) or Adam via trivializations (Lezcano Casado,, 2019)) could further reduce this dependence.
@@ E.1 Performance Across Data Distributions
NJ HS RHS (ours) vs HS (RED) vs NJ (RED)
$|P|$ RED $(\uparrow)$ CPU RED $(\uparrow)$ CPU RED $(\uparrow)$ CPU $p_{\text{Welch}}$ $p_{\text{TOST}}$ $p_{\text{Welch}}$ $p_{\text{TOST}}$
10 $1.46_{\pm 1.12}$ $5.50_{\pm 0.03}$ $2.72_{\pm 0.97}$ $0.03_{\pm 0.01}$ $3.17_{\pm 1.59}$ $4.60_{\pm 3.40}$ .457.457 .183.183 $\mathbf{.013}$ .867.867
20 $2.60_{\pm 3.00}$ $5.68_{\pm 0.04}$ $1.83_{\pm 0.70}$ $0.06_{\pm 0.01}$ $3.15_{\pm 0.56}$ $9.87_{\pm 7.00}$ $\mathbf{<\!.001}$ .863.863 $\mathbf{<\!.001}$ $1.000$
30 $1.70_{\pm 2.31}$ $5.86_{\pm 0.03}$ $2.12_{\pm 0.60}$ $0.10_{\pm 0.02}$ $3.03_{\pm 1.09}$ $12.65_{\pm 6.94}$ $\mathbf{.036}$ .411.411 $\mathbf{<\!.001}$ $1.000$
40 $2.62_{\pm 1.35}$ $5.95_{\pm 0.04}$ $2.17_{\pm 0.71}$ $0.13_{\pm 0.02}$ $3.30_{\pm 0.75}$ $17.59_{\pm 6.97}$ $\mathbf{.003}$ .652.652 $\mathbf{<\!.001}$ $1.000$
50 $4.37_{\pm 1.87}$ $6.14_{\pm 0.04}$ $2.26_{\pm 0.43}$ $0.17_{\pm 0.02}$ $3.08_{\pm 0.34}$ $26.21_{\pm 8.67}$ $\mathbf{<\!.001}$ .157.157 $\mathbf{<\!.001}$ $1.000$
60 $3.38_{\pm 1.09}$ $6.29_{\pm 0.03}$ $2.56_{\pm 0.57}$ $0.20_{\pm 0.02}$ $2.72_{\pm 0.48}$ $26.95_{\pm 8.40}$ .506.506 $\mathbf{.001}$ $\mathbf{<\!.001}$ $1.000$
70 $3.28_{\pm 1.48}$ $6.51_{\pm 0.04}$ $2.45_{\pm 0.54}$ $0.22_{\pm 0.01}$ $2.68_{\pm 0.66}$ $31.09_{\pm 11.84}$ .405.405 $\mathbf{.005}$ $\mathbf{<\!.001}$ $1.000$
80 $3.89_{\pm 1.79}$ $6.77_{\pm 0.03}$ $2.43_{\pm 0.53}$ $0.26_{\pm 0.02}$ $3.07_{\pm 0.39}$ $38.55_{\pm 8.45}$ $\mathbf{.007}$ .051.051 $\mathbf{<\!.001}$ $1.000$
90 $3.46_{\pm 1.65}$ $7.01_{\pm 0.04}$ $2.72_{\pm 0.50}$ $0.30_{\pm 0.02}$ $2.87_{\pm 0.57}$ $66.07_{\pm 30.15}$ .540.540 $\mathbf{.001}$ $\mathbf{<\!.001}$ $1.000$
100 $3.77_{\pm 1.75}$ $7.24_{\pm 0.03}$ $2.45_{\pm 0.35}$ $0.33_{\pm 0.02}$ $2.84_{\pm 0.47}$ $70.31_{\pm 14.92}$ .051.051 $\mathbf{.002}$ $\mathbf{<\!.001}$ $1.000$
150 $4.17_{\pm 1.10}$ $8.46_{\pm 0.03}$ $2.27_{\pm 0.20}$ $0.48_{\pm 0.02}$ $2.48_{\pm 0.22}$ $139.87_{\pm 49.03}$ $\mathbf{.039}$ $\mathbf{<\!.001}$ $\mathbf{<\!.001}$ $1.000$
200 $5.01_{\pm 1.40}$ $11.09_{\pm 0.06}$ $2.43_{\pm 0.26}$ $0.66_{\pm 0.04}$ $2.64_{\pm 0.23}$ $341.42_{\pm 96.02}$ .072.072 $\mathbf{<\!.001}$ $\mathbf{<\!.001}$ $1.000$
@@
Table 1: Sampling $|P|$ points from a centered hyperbolic Gaussian $\mathcal{G}(0,0.5)$. RED: reduction over MST (%). CPU: total CPU time (sec.). Significance columns report Welch $t$-test and TOST p-values (RHS vs. the indicated baseline, RED metric, n=10n{=}10, $\varepsilon{=}1.0$). Bold indicates p<0.05p{<}0.05; marks TOST equivalence (90 % CI $\subseteq(-\varepsilon,+\varepsilon)$ and $p_{\text{TOST}}{<}0.05$).
NJ HS RHS (ours) vs HS (RED) vs NJ (RED)
$|P|$ RED $(\uparrow)$ CPU RED $(\uparrow)$ CPU RED $(\uparrow)$ CPU $p_{\text{Welch}}$ $p_{\text{TOST}}$ $p_{\text{Welch}}$ $p_{\text{TOST}}$
10 $40.49_{\pm 0.03}$ $5.52_{\pm 0.04}$ $17.75_{\pm 4.57}$ $0.02_{\pm 0.01}$ $28.33_{\pm 18.54}$ $25.70_{\pm 13.30}$ .110.110 .928.928 .068.068 .955.955
20 $40.10_{\pm 0.06}$ $5.73_{\pm 0.05}$ $14.35_{\pm 4.61}$ $0.05_{\pm 0.02}$ $40.01_{\pm 0.09}$ $61.94_{\pm 16.77}$ $\mathbf{<\!.001}$ $1.000$ $\mathbf{.018}$ $\mathbf{<\!.001}$
30 $39.93_{\pm 0.08}$ $5.91_{\pm 0.04}$ $14.74_{\pm 5.75}$ $0.08_{\pm 0.06}$ $39.84_{\pm 0.13}$ $95.05_{\pm 27.50}$ $\mathbf{<\!.001}$ $1.000$ .082.082 $\mathbf{<\!.001}$
40 $39.70_{\pm 0.09}$ $6.06_{\pm 0.03}$ $16.63_{\pm 4.46}$ $0.09_{\pm 0.03}$ $39.71_{\pm 0.12}$ $119.86_{\pm 53.15}$ $\mathbf{<\!.001}$ $1.000$ .836.836 $\mathbf{<\!.001}$
50 $39.53_{\pm 0.06}$ $6.00_{\pm 0.02}$ $17.88_{\pm 5.29}$ $0.10_{\pm 0.04}$ $39.52_{\pm 0.05}$ $116.38_{\pm 45.92}$ $\mathbf{<\!.001}$ $1.000$ .690.690 $\mathbf{<\!.001}$
60 $39.22_{\pm 0.10}$ $6.37_{\pm 0.04}$ $14.95_{\pm 4.60}$ $0.11_{\pm 0.02}$ $39.36_{\pm 0.11}$ $87.99_{\pm 9.33}$ $\mathbf{<\!.001}$ $1.000$ $\mathbf{.008}$ $\mathbf{<\!.001}$
70 $39.01_{\pm 0.16}$ $6.55_{\pm 0.05}$ $14.42_{\pm 4.27}$ $0.11_{\pm 0.03}$ $39.26_{\pm 0.05}$ $118.26_{\pm 24.23}$ $\mathbf{<\!.001}$ $1.000$ $\mathbf{<\!.001}$ $\mathbf{<\!.001}$
80 $38.55_{\pm 0.11}$ $6.77_{\pm 0.04}$ $14.47_{\pm 4.29}$ $0.12_{\pm 0.03}$ $39.11_{\pm 0.14}$ $111.56_{\pm 12.72}$ $\mathbf{<\!.001}$ $1.000$ $\mathbf{<\!.001}$ $\mathbf{<\!.001}$
90 $38.18_{\pm 0.18}$ $6.96_{\pm 0.06}$ $15.85_{\pm 3.70}$ $0.14_{\pm 0.03}$ $39.03_{\pm 0.08}$ $171.84_{\pm 59.68}$ $\mathbf{<\!.001}$ $1.000$ $\mathbf{<\!.001}$ $\mathbf{.016}$
100 $37.50_{\pm 0.28}$ $7.26_{\pm 0.07}$ $13.91_{\pm 4.99}$ $0.13_{\pm 0.04}$ $38.94_{\pm 0.12}$ $183.57_{\pm 60.30}$ $\mathbf{<\!.001}$ $1.000$ $\mathbf{<\!.001}$ $1.000$
150 $24.02_{\pm 1.24}$ $8.48_{\pm 0.05}$ $13.22_{\pm 6.21}$ $0.16_{\pm 0.02}$ $38.47_{\pm 0.10}$ $205.34_{\pm 52.63}$ $\mathbf{<\!.001}$ $1.000$ $\mathbf{<\!.001}$ $1.000$
200 $8.08_{\pm 1.09}$ $10.51_{\pm 0.06}$ $13.62_{\pm 4.43}$ $0.18_{\pm 0.03}$ $38.17_{\pm 0.17}$ $265.46_{\pm 80.59}$ $\mathbf{<\!.001}$ $1.000$ $\mathbf{<\!.001}$ $1.000$
Table 2: Sampling $|P|$ points from a mixture of $10$ hyperbolic Gaussians $\mathcal{G}(\mu_{10,k}(1-10^{-10}),0.1)$, $k\in\{1,\ldots,10\}$. RED: reduction over MST (%). CPU: total CPU time (sec.). Significance columns report Welch $t$-test and TOST p-values (RHS vs. the indicated baseline, RED metric, n=10n{=}10, $\varepsilon{=}1.0$). Bold indicates p<0.05p{<}0.05; marks TOST equivalence (90 % CI $\subseteq(-\varepsilon,+\varepsilon)$ and $p_{\text{TOST}}{<}0.05$).
For each configuration, we assess statistical differences between RHS and each baseline via two complementary tests: the Welch–Satterthwaite two-sample $t$-test (Welch,, 1938; Satterthwaite,, 1946), which detects significant differences in means without assuming equal variances across groups, and the Two One-Sided Tests (TOST) procedure (Schuirmann,, 1987), which declares practical equivalence when the 90% confidence interval of the mean difference lies within $\varepsilon,+\varepsilon]$ (in our case we choose $\varepsilon{=}1.0$). A significant Welch $p$-value with a non-significant TOST indicates a detectable difference; a significant TOST () with a non-significant Welch confirms practical equivalence; both significant implies a real but negligible gap; neither significant is inconclusive.